OpenStudy (anonymous):

Evaluate the integral S(sigma) (1/√x × e^√x) dx

OpenStudy (anonymous):

1/sqrt(x e^sqrt(x)) or 1/(sqrt(x)e^sqrt(x))

OpenStudy (anonymous):

which function are you trying to integrate

OpenStudy (anonymous):

\[\int\limits_{}^{} (1/\sqrt{x} \times e ^{\sqrt{x}} ) dx\]

OpenStudy (anonymous):

im not sure which function its just a problem a have from a worksheet

OpenStudy (anonymous):

u=sqrt(x)

OpenStudy (anonymous):

try u substitution

OpenStudy (anonymous):

am i substituting the exponent sqrt of the base sqrt??

OpenStudy (anonymous):

*or

OpenStudy (anonymous):

substitute the exponent sqrt

OpenStudy (anonymous):

you can also substitute both if you want to

OpenStudy (anonymous):

ill do the exponent caould it be 1/2x^-1/2 or do i have to do U^2=...

OpenStudy (anonymous):

du = 1/2 x^(-1/2) dx works fine

OpenStudy (anonymous):

the x^(-1/2) will cancel with base sqrt, when you convert dx to du

OpenStudy (anonymous):

so so far its 1/2e^sqrt(u)

OpenStudy (anonymous):

u=sqrt(x) du = 1/2 x^(-1/2) dx 1/(sqrt(x)e^sqrt(x))dx=1/(sqrt(x)e^u)dx 1/(sqrt(x)e^u)dx = 2sqrt(x)/(sqrt(x)e^u) du

OpenStudy (anonymous):

du = 1/2 x^(-1/2) dx dx = 2sqrt(x)du

OpenStudy (anonymous):

okay i understand wat i did different

OpenStudy (anonymous):

what will i need to do next??

OpenStudy (anonymous):

1/(sqrt(x)e^u)dx = 2sqrt(x)/(sqrt(x)e^u) du 2sqrt(x)/(sqrt(x)e^u) du = 2/e^u du integrate this, then replace u with sqrt(x) after integrating

OpenStudy (anonymous):

so i would need to put ln(2/e^sqrt(x)) +c

OpenStudy (anonymous):

how do i change 2/e^sqrtx

OpenStudy (anonymous):

where is the ln coming from?

OpenStudy (anonymous):

just integrate 2/e^u du =2 e^(-u) du =-2e^(-u)+C then replace u with sqrt(x) =-2e^(-sqrt(x))+C

OpenStudy (anonymous):

hmm my teacher would always use ln wen integrating

OpenStudy (anonymous):

but i get how you do it

OpenStudy (anonymous):

why did you make 2 negative if you only brought up e^sqrtx

OpenStudy (anonymous):

integrate(2 e^(-u) du)=-2e^(-u)+C the 2 becomes negative after taking the integral integrate(e^(-x)) = -e^(-x)

OpenStudy (anonymous):

okay thnks for your help i really appreciate it